# Finite presentability of twisted Brin–Thompson groups

### v1

@@ 3. Finite presentability from actions with bad edge stabilizers
Let us say triangle for $2$-simplex. From each orbit of triangles in $K^{(2)}$, we can choose one that has at least one vertex in $T$. Given a triangle $\Delta$ in $K^{(2)}$ with a vertex $v$ in $T$, view the boundary of $\Delta$ as an edge path (of length three) from $v$ to $v$, say $e_{1},e_{2},e_{3}$. (Up to possibly switching the roles of some $e_{i}$ and $\overline{e}_{i}$, we can assume that this is a directed cycle, i.e., $t(e_{i})=o(e_{i+1})$ for all $i$ mod $3$.) For each $i=1,2,3$, let $f_{i}$ be the image of $e_{i}$ in $\Gamma\backslash K^{(1)}$, and let $\widetilde{f}_{i}$ be the lift as above. If $f_{i}$ is in $M$ then $\widetilde{f}_{i}$ is in $T$, and if not then the origin $o(\widetilde{f}_{i})$ lies in $T$ but not the terminus. Since $e_{1}$ and $\widetilde{f}_{1}$ share a $\Gamma$-orbit, and both of their origins lie in $T$, they must have the same origin, namely $v$. Since they share an orbit and have the same origin $v$, and the action is orientation preserving, we can choose $a_{1}\in\Gamma_{v}\coloneqq\operatorname{Stab}_{\Gamma}(v)$ taking $\widetilde{f}_{1}$ to $e_{1}$. Now observe that $a_{1}\gamma_{f_{1}}$ takes $z_{f_{1}}$ to the terminus of $e_{1}$, which is the origin of $e_{2}$. Since the origin of $\widetilde{f}_{2}$ is the vertex of $T$ sharing an orbit with the origin of $e_{2}$, the origin of $\widetilde{f}_{2}$ is $z_{f_{1}}$. Now similar to before, we can choose $a_{2}\in\Gamma_{z_{f_{1}}}$ such that $a_{1}\gamma_{f_{1}}a_{2}$ takes $\widetilde{f}_{2}$ to $e_{2}$, and so $a_{1}\gamma_{f_{1}}a_{2}\gamma_{f_{2}}$ takes $z_{f_{2}}$ to the terminus of $e_{2}$, which is the origin of $e_{3}$. Finally, we do this trick again and get $a_{3}\in\Gamma_{z_{f_{2}}}$ such that $a_{1}\gamma_{f_{1}}a_{2}\gamma_{f_{2}}a_{3}\gamma_{f_{3}}$ takes $z_{f_{3}}$ to the terminus of $e_{3}$, which is $v$. But $z_{f_{3}}$ is a vertex in $T$ sharing an orbit with $v$, hence must be $v$. We conclude that $a_{1}\gamma_{f_{1}}a_{2}\gamma_{f_{2}}a_{3}\gamma_{f_{3}}$ lies in $\Gamma_{v}$, and so up to modifying $a_{1}$ we have $a_{1}\gamma_{f_{1}}a_{2}\gamma_{f_{2}}a_{3}\gamma_{f_{3}}=1$. Call this relation $r_{\Delta}$. Note that we can phrase $r_{\Delta}$ as saying that $a_{1}\gamma_{f_{1}}a_{2}\gamma_{f_{2}}a_{3}=\gamma_{f_{3}}^{-1}$, so in particular if we have already chosen $\gamma_{f_{1}}$ and $\gamma_{f_{2}}$, then we can choose $\gamma_{f_{3}}$ in terms of $\gamma_{f_{1}}$, $\gamma_{f_{2}}$, and elements of the various vertex stabilizers.
@@
for all $f$ in $(\Gamma\backslash K^{(1)})\setminus M$ and $g\in\Gamma_{\widetilde{f}}$, called edge relations. (Here as usual $x^{y}\coloneqq y^{-1}xy$.) For a fixed $f$, write $R(f,*)$ for the family of such relations. Note that all the $R(f,g)$ are consequences of those for which $g$ comes from a generating set of $\Gamma_{\widetilde{f}}$. At this point it is clear that if each $\Gamma_{v}$ is finitely presented and each $\Gamma_{\widetilde{f}}$ is finitely generated, and the action is cocompact, then $\Gamma$ is finitely presented. Now we will inspect the case when not every $\Gamma_{\widetilde{f}}$ is finitely generated.
@@
Let $f$ be an edge in $\Gamma\backslash K^{(1)}$ such that $\Gamma_{\widetilde{f}}$ is not finitely generated. Let $e_{1},\dots,e_{n}$ be an edge path in $K^{(1)}$ from $v=o(\widetilde{f})$ to $w=t(\widetilde{f})$ such that each $\Gamma_{e_{i}}$ is finitely generated, and $\bigcap\limits_{i=1}^{n}\Gamma_{e_{i}}$ has finite index in $\Gamma_{\widetilde{f}}$. Let $f_{i}$ be the image of $e_{i}$ in $\Gamma\backslash K^{(1)}$, with $\widetilde{f}_{i}$ the chosen lift to $K^{(1)}$ (so $o(\widetilde{f}_{i})$ lies in $T$). We now claim that the edge relations $R(f,*)$ are all consequences of the edge relations $R(f_{i},*)$ together with the (finitely many) non-edge relations. By extrapolating the triangle relations, and noting that the path $e_{1},\dots,e_{n},\overline{f}$ is a directed cycle, we see that we are free to choose $\gamma_{f}=\gamma_{\overline{f}}^{-1}$ so that
@@ 4. Stein complexes and subcomplexes
Next, define a multicolored tree recursively, by saying that the identity on $\mathfrak{C}^{S}$ and all simple splits $x_{s}$ are multicolored trees, and if $f_{1}$ and $f_{2}$ are multicolored trees and $\sigma$ is a permutation, then $f=p_{\sigma}(f_{1}\oplus f_{2})x_{s}$ is a multicolored tree. The name comes from viewing $S$ as a set of colors, and $x_{s}$ as a caret with color $s$. A multicolored forest is any element of $S\mathcal{V}$ of the form $p_{\sigma}(f_{1}\oplus\cdots\oplus f_{m})$ for multicolored trees $f_{i}$ and $\sigma$ a permutation. Note that the set of all multicolored forests is the smallest subset of $S\mathcal{V}$ that contains all permutations and simple splits, and is closed under direct sums and compositions.
Let $P$ be the set of equivalence classes $[h]=\mathcal{G}(m)h$, where $h$ is an element of $S\mathcal{V}_{G}$ with rank $m$. That is, $P$ is the set of elements of the twisted Brin–Thompson groupoid, up to postcomposing by twisted permutations. Let $P_{1}$ be the subposet of elements with corank $1$. We have a (right) action of $SV_{G}$ on $P_{1}$ given by precomposing. The set $P$ has a partial order $\leq$, given by
@@
whenever $f$ is a multicolored forest (whose corank equals the rank of $h$). By [BZ22, Lemma 5.1 and Proposition 5.2], $\leq$ is a partial order and the subposet $P_{1}$ is directed, so the geometric realization $|P_{1}|$ is contractible. One key to $\leq$ being a partial order is the following result, which will also be important here.
@@
For $[h]\preceq[fh]$ in $P_{1}$, write $[h]\preceq_{k}[fh]$ whenever the multicolored forest $f$ is $k$-elementary. Given a simplex in $|P_{1}|$, i.e., a chain $v_{0}<v_{1}<\cdots<v_{k}$ of elements of $P_{1}$, call the simplex $k$-elementary if $v_{0}\preceq_{k}v_{k}$ (and hence $v_{i}\preceq_{k}v_{j}$ for all $i<j$). These simplices form a subcomplex of $|P_{1}|$, which we denote by $X(k)$, and call the $k$-elementary Stein complex for $SV_{G}$. The Stein complex $X$ is $X=X(\infty)$.
@@
By [BZ22, Proposition 7.9], which establishes connectivity properties of the descending links, together with standard discrete Morse theory, e.g., [BB97, Corollary 2.6], one can compute that for any $n$, the complex $X_{m}$ is $(n-1)$-connected for $m$ sufficiently large. For example, $X_{m}$ is simply connected for all $m\geq 15$.
@@
Let $k\geq n$. The complex $X(k)$ is $(n-1)$-connected, and moreover for any $m\in\mathbb{N}$, if $m$ is large enough that the truncation $X_{m}$ is $(n-1)$-connected then $X_{m}(k)$ is also $(n-1)$-connected. For example, $X_{15}(2)$ is simply connected.
@@
Say $v=[h]$ and $w=[fh]$ for $f$ an elementary (and not $k$-elementary) multicolored forest. Let $r$ be the rank of $h$, which is also the corank of $f$. For any $u\in(v,w)$, since $v<u$ we can write $u=[f^{\prime}h]$ for some non-trivial multicolored forest $f^{\prime}$ (here we have used Lemma 4.1), so in particular there exists a simple split $x_{s}$ such that $v<[(\operatorname{id}_{p}\oplus x_{s}\oplus\operatorname{id}_{r-p-1})h]\leq u<w$ for some $10\leq p\leq r-$ 1. Here $\operatorname{id}_{p}$ is the identity element with rank (and corank) $p$. In particular, every simplex of $|(v,w)|$ lies in the star of a vertex of $|(v,w)|$ of the form $u_{p}(s)\coloneqq[(\operatorname{id}_{p}\oplus x_{s}\oplus\operatorname{id}_{r-p-1})h]$. The complex $|(v,w)|$ is therefore covered by these stars, and we can use a Nerve Lemma, for example [BLVŽ94, Lemma 1.2], to inspect its connectivity.
@@
The action of $SV_{G}$ on $|P_{1}|$ stabilizes $X_{m}(k)$ for each $m$ and $k$, so we can inspect the action of $SV_{G}$ on $X_{m}(k)$. If $m$ and $k$ are large enough, this is simply connected, for example $X_{15}(2)$. In order to get $SV_{G}$ to be finitely presented, we need to know things about cocompactness and about stabilizers. First we establish cocompactness
@@
We can also establish some results about stabilizers. Recall that two groups are commensurable if they have isomorphic finite index subgroups. Note that any group commensurable to a group of type $\operatorname{F}_{n}$ is also of type $\operatorname{F}_{n}$, and a direct product of finitely many groups of type $\operatorname{F}_{n}$ is of type $\operatorname{F}_{n}$.
@@
The forward direction is Proposition 2.1. Now we do the backward direction. Suppose $G$ is a group with an action of type (A) on a set $S$, and we need to prove that $SV_{G}$ is finitely presented. Consider the (orientation preserving) action of $SV_{G}$ on the complex $X_{m}(k)$, for $m\geq 15$ and $k\geq 2$. This complex is simply connected by Proposition 4.4, and the action is cocompact by Lemma 4.5. Every vertex stabilizer is finitely presented by Citation 4.6. Any stabilizer of a short edge is finitely generated by Citation 4.8. Finally, for any long edge, by Lemma 4.9 there is a path of short edges connecting its endpoints, such that the fixer of the path has finite index in the stabilizer of the long edge. Now Proposition 3.1 tells us that $SV_{G}$ is finitely presented. ∎
@@ 5. Applications to the Boone–Higman conjecture
Recall Corollary B, that any subgroup of a group admitting an action of type (A) satisfies the Boone–Higman conjecture. In this section, we investigate this further.
5.1. Boone–Higman via (A)
First, let us define the following sufficient condition for satisfying the Boone–Higman conjecture:
Definition 5.1 (Boone–Higman via (A)).
We say that a group satisfies the Boone–Higman conjecture via (A) if it embeds into a group admitting an action of type (A).
We should mention that we do not know whether a group can satisfy the Boone–Higman conjecture without doing so via (A). Let us also emphasize that when we say a group satisfies the Boone–Higman conjecture via (A), we do not require the group itself to admit an action of type (A), just that it embeds into a group that does.
Let us give a new example of a family of groups satisfying the Boone–Higman conjecture via (A). Our examples come from so called shift-similar groups, introduced by Mallery and the author in [MZ]. These can be viewed as an analog of self-similar groups (as in [Nek05]), that relate to Houghton groups rather than Thompson groups.

### v2

@@ 3. Finite presentability from actions with bad edge stabilizers
Let us say triangle for $2$-simplex. From each orbit of triangles in $K^{(2)}$, we can choose one that has at least one vertex in $T$. Given a triangle $\Delta$ in $K^{(2)}$ with a vertex $v$ in $T$, view the boundary of $\Delta$ as an edge path (of length three) from $v$ to $v$, say $e_{1},e_{2},e_{3}$. (Up to possibly switching the roles of some $e_{i}$ and $\overline{e}_{i}$, we can assume that this is a directed cycle, i.e., $t(e_{i})=o(e_{i+1})$ for all $i$ mod $3$.) For each $i=1,2,3$, let $f_{i}$ be the image of $e_{i}$ in $\Gamma\backslash K^{(1)}$, and let $\widetilde{f}_{i}$ be the lift as above. If $f_{i}$ is in $M$ then $\widetilde{f}_{i}$ is in $T$, and if not then the origin $o(\widetilde{f}_{i})$ lies in $T$ but not the terminus. Since $e_{1}$ and $\widetilde{f}_{1}$ share a $\Gamma$-orbit, and both of their origins lie in $T$, they must have the same origin, namely $v$. Since they share an orbit and have the same origin $v$, and the action is orientation preserving, we can choose $a_{1}\in\Gamma_{v}\coloneqq\operatorname{Stab}_{\Gamma}(v)$ taking $\widetilde{f}_{1}$ to $e_{1}$. Now observe that $a_{1}\gamma_{f_{1}}$ takes $z_{f_{1}}$ to the terminus of $e_{1}$, which is the origin of $e_{2}$. Since the origin of $\widetilde{f}_{2}$ is the vertex of $T$ sharing an orbit with the origin of $e_{2}$, the origin of $\widetilde{f}_{2}$ is $z_{f_{1}}$. Now similar to before, we can choose $a_{2}\in\Gamma_{z_{f_{1}}}$ such that $a_{1}\gamma_{f_{1}}a_{2}$ takes $\widetilde{f}_{2}$ to $e_{2}$, and so $a_{1}\gamma_{f_{1}}a_{2}\gamma_{f_{2}}$ takes $z_{f_{2}}$ to the terminus of $e_{2}$, which is the origin of $e_{3}$. Finally, we do this trick again and get $a_{3}\in\Gamma_{z_{f_{2}}}$ such that $a_{1}\gamma_{f_{1}}a_{2}\gamma_{f_{2}}a_{3}\gamma_{f_{3}}$ takes $z_{f_{3}}$ to the terminus of $e_{3}$, which is $v$. But $z_{f_{3}}$ is a vertex in $T$ sharing an orbit with $v$, hence must be $v$. We conclude that $a_{1}\gamma_{f_{1}}a_{2}\gamma_{f_{2}}a_{3}\gamma_{f_{3}}$ lies in $\Gamma_{v}$, write $a_{1}\gamma_{f_{1}}a_{2}\gamma_{f_{2}}a_{3}\gamma_{f_{3}}=a_{v}$. Now we have $a_{v}^{-1}a_{1}\gamma_{f_{1}}a_{2}\gamma_{f_{2}}a_{3}\gamma_{f_{3}}=1$, and up to forgetting how $a_{1}$ arose in $\Gamma_{v}$ specifically taking $\widetilde{f}_{1}$ to $e_{1}$ we can rename $a_{v}^{-1}a_{1}$ as the new $a_{1}$. Thus we have a relation
$a_{1}\gamma_{f_{1}}a_{2}\gamma_{f_{2}}a_{3}\gamma_{f_{3}}=1\text{,}$
with $a_{1}$, $a_{2}$, and $a_{3}$ coming from various vertex stabilizers, and we call this relation $r_{\Delta}$. We can phrase $r_{\Delta}$ as saying that $a_{1}\gamma_{f_{1}}a_{2}\gamma_{f_{2}}a_{3}=\gamma_{f_{3}}^{-1}$, so in particular if we have already chosen $\gamma_{f_{1}}$ and $\gamma_{f_{2}}$, then we can choose $\gamma_{f_{3}}$ in terms of $\gamma_{f_{1}}$, $\gamma_{f_{2}}$, and elements of the various vertex stabilizers.
@@
for all $f$ in $(\Gamma\backslash K^{(1)})\setminus M$ and $g\in\Gamma_{\widetilde{f}}$, called edge relations. Here as usual $x^{y}\coloneqq y^{-1}xy$. (To be more clear about $\gamma$ versus $\lambda$, the $\gamma_{f}$ are elements of the group $\Gamma$, whereas the $\lambda_{f}$ are formal symbols used as generators in the abstract group presentation. The tree relations $\lambda_{f}=1$ essentially account for the fact that we chose $\gamma_{f}=1$ whenever $f$ in the tree $M$.)
For a fixed $f$, write $R(f,*)$ for the family of relations $R(f,g)$ for $g\in\Gamma_{\widetilde{f}}$. Note that all the $R(f,g)$ are consequences of those for which $g$ comes from a generating set of $\Gamma_{\widetilde{f}}$. At this point it is clear that if each $\Gamma_{v}$ is finitely presented and each $\Gamma_{\widetilde{f}}$ is finitely generated, and the action is cocompact, then $\Gamma$ is finitely presented. Now we will inspect the case when not every $\Gamma_{\widetilde{f}}$ is finitely generated.
@@
Let $f$ be an edge in $\Gamma\backslash K^{(1)}$ such that $\Gamma_{\widetilde{f}}$ is not finitely generated. Let $e_{1},\dots,e_{n}$ be an edge path in $K^{(1)}$ from $v=o(\widetilde{f})$ to $w=t(\widetilde{f})$ such that each $\Gamma_{e_{i}}$ is finitely generated, and $\bigcap\limits_{i=1}^{n}\Gamma_{e_{i}}$ has finite index in $\Gamma_{\widetilde{f}}$. Let $f_{i}$ be the image of $e_{i}$ in $\Gamma\backslash K^{(1)}$, with $\widetilde{f}_{i}$ the chosen lift to $K^{(1)}$ (so $o(\widetilde{f}_{i})$ lies in $T$). We now claim that the edge relations $R(f,*)$ are all consequences of the edge relations $R(f_{i},*)$ together with the (finitely many) non-edge relations. By extrapolating the triangle relations, and noting that the path $e_{1},\dots,e_{n},\overline{f}$ is a directed cycle, we see that we are free to choose $\gamma_{f}$ (which equals $\gamma_{\overline{f}}^{-1}$) so that
@@ 4. Stein complexes and subcomplexes
Next, define a multicolored tree recursively, by saying that the identity on $\mathfrak{C}^{S}$ and all simple splits $x_{s}$ are multicolored trees, and if $f_{1}$ and $f_{2}$ are multicolored trees and $\sigma$ is a permutation, then $f=p_{\sigma}(f_{1}\oplus f_{2})x_{s}$ is a multicolored tree. The name comes from viewing $S$ as a set of colors, and $x_{s}$ as a caret with color $s$. Note that multicolored trees have corank $1$, and can have any rank. A multicolored forest is any element of $S\mathcal{V}$ of the form $p_{\sigma}(f_{1}\oplus\cdots\oplus f_{m})$ for multicolored trees $f_{i}$ and $\sigma$ a permutation. Note that the set of all multicolored forests is the smallest subset of $S\mathcal{V}$ that contains all permutations and simple splits, and is closed under direct sums and compositions.
Let $P$ be the set of equivalence classes $[h]=\mathcal{G}(m)h$, where $h$ is an element of $S\mathcal{V}_{G}$ with rank $m$. That is, $P$ is the set of elements of the twisted Brin–Thompson groupoid, up to postcomposing by twisted permutations. Let $P_{1}$ be the subset of elements with corank $1$. We have a (right) action of $SV_{G}$ on $P_{1}$ given by precomposing. The set $P$ has a partial order $\leq$, given by
@@
whenever $f$ is a multicolored forest (whose corank equals the rank of $h$). Note that $P_{1}$ is a subposet of $P$. By [BZ22, Lemma 5.1 and Proposition 5.2], $\leq$ is a partial order and the subposet $P_{1}$ is directed, so the geometric realization $|P_{1}|$ is contractible. One key to $\leq$ being a partial order is the following result, which will also be important here.
@@
For $[h]\preceq[fh]$ in $P_{1}$, write $[h]\preceq_{k}[fh]$ whenever the multicolored forest $f$ is $k$-elementary. Given a simplex in $|P_{1}|$, i.e., a chain $v_{0}<v_{1}<\cdots<v_{k}$ of elements of $P_{1}$, call the simplex $k$-elementary if $v_{0}\preceq_{k}v_{k}$ (and hence $v_{i}\preceq_{k}v_{j}$ for all $i<j$). These simplices form a subcomplex of $|P_{1}|$, which we denote by $X(k)$, and call the $k$-elementary Stein complex for $SV_{G}$. The Stein complex $X$ is $X=X(\infty)$, i.e., the complex dictated by using all elementary multicolored forests rather than restricting to just the $k$-elementary ones for some $k$.
@@
By [BZ22, Proposition 7.9], which establishes connectivity properties of the descending links, together with standard discrete Morse theory, e.g., [BB97, Corollary 2.6], one can compute that for any $n$, the complex $X_{m}$ is $(n-1)$-connected for $m$ sufficiently large. For example, $X_{m}$ is simply connected for all $m\geq 21$, since [BZ22, Proposition 7.9] says the descending link of any vertex with rank greater than $m$ is simply connected as soon as $\lfloor((m+1)/2-2)/3\rfloor-2\geq 1$ and $\log_{2}((m+1)/2)-2\geq 1$, which is equivalent to $m\geq 21$.
@@
Let $k\geq n$. The complex $X(k)$ is $(n-1)$-connected, and moreover for any $m\in\mathbb{N}$, if $m$ is large enough that the truncation $X_{m}$ is $(n-1)$-connected then $X_{m}(k)$ is also $(n-1)$-connected. For example, $X_{21}(2)$ is simply connected.
@@
Say $v=[h]$ and $w=[fh]$ for $f$ an elementary (and not $k$-elementary) multicolored forest. Let $r$ be the rank of $h$, which is also the corank of $f$. For any $u\in(v,w)$, since $v<u$ we can write $u=[f^{\prime}h]$ for some non-trivial multicolored forest $f^{\prime}$ (here we have used Lemma 4.1), so in particular there exists a simple split $x_{s}$ such that $v<[(\operatorname{id}_{p}\oplus x_{s}\oplus\operatorname{id}_{r-p-1})h]\leq u<w$ for some $10\leq p\leq r-$ 1. Here $\operatorname{id}_{p}$ is the identity element with rank (and corank) $p$. In particular, every simplex of $|(v,w)|$ lies in the star of a vertex of $|(v,w)|$ of the form $u_{p}(s)\coloneqq[(\operatorname{id}_{p}\oplus x_{s}\oplus\operatorname{id}_{r-p-1})h]$. The complex $|(v,w)|$ is therefore covered by these stars, and we can use a Nerve Lemma, for example [BLVŽ94, Lemma 1.2], to inspect its connectivity. Note that the distinguishing feature of these vertices $u_{p}(s)$ is that $v\preceq_{1}u_{p}(s)$.
@@
The action of $SV_{G}$ on $|P_{1}|$ stabilizes $X_{m}(k)$ for each $m$ and $k$, so we can inspect the action of $SV_{G}$ on $X_{m}(k)$. If $m$ and $k$ are large enough, this is simply connected, for example $X_{21}(2)$. In order to get $SV_{G}$ to be finitely presented, we need to know things about cocompactness and about stabilizers. First we establish cocompactness
@@
We can also establish some results about stabilizers. Recall that two groups are commensurable if they have isomorphic finite index subgroups. Note that any group commensurable to a group of type $\operatorname{F}_{n}$ is also of type $\operatorname{F}_{n}$, and a direct product of finitely many groups of type $\operatorname{F}_{n}$ is of type $\operatorname{F}_{n}$. This is because, on the level of classifying spaces, a complex has finite $n$-skeleton if and only if every finite-sheeted cover does, and a product of complexes with finite $n$-skeleta has finite $n$-skeleton.
@@
The forward direction is Proposition 2.1. Now we do the backward direction. Suppose $G$ is a group with an action of type (A) on a set $S$, and we need to prove that $SV_{G}$ is finitely presented. Consider the (orientation preserving) action of $SV_{G}$ on the complex $X_{m}(2)$, for $m\geq 21$. This complex is simply connected by Proposition 4.4, and the action is cocompact by Lemma 4.5. Every vertex stabilizer is finitely presented by Citation 4.6. Any stabilizer of a short edge is finitely generated by Citation 4.8. Finally, for any long edge, by Lemma 4.9 there is a path of short edges connecting its endpoints, such that the fixer of the path has finite index in the stabilizer of the long edge. Now Proposition 3.1 tells us that $SV_{G}$ is finitely presented. ∎
@@ 5. Applications to the Boone–Higman conjecture
Recall Corollary B, that any subgroup of a group admitting an action of type (A) (has solvable word problem and) satisfies the Boone–Higman conjecture. In this section, we investigate this further.
5.1. Permutational Boone–Higman
First, let us define the following relative of the Boone–Higman conjecture:
Conjecture 5.1 (Permutational Boone–Higman conjecture).
A finitely generated group has solvable word problem if and only if it embeds in a group admitting an action of type (A).
The “if” direction is clear from the “if” direction of the Boone–Higman conjecture together with Theorem A. When we say a group “satisfies” this conjecture, we mean it (has solvable word problem and) embeds in a group admitting an action of type (A). We do not know whether a group can satisfy the Boone–Higman conjecture without satisfying the permutational version (see Question 5.13). Let us also emphasize that when we say a group satisfies the permutational Boone–Higman conjecture, we do not require the group itself to admit an action of type (A), just that it embeds into a group that does.
Let us give a new example of a family of groups satisfying the permutational Boone–Higman conjecture, and hence the Boone–Higman conjecture. Our examples come from so called shift-similar groups, introduced by Mallery and the author in [MZ]. These can be viewed as an analog of self-similar groups (as in [Nek05]), that relate to Houghton groups rather than Thompson groups.